OpenStudy (anonymous):

22x+70=17x-95 is my equation and after all the work i got x= 11 but when i checked it ,it wasn't balanced here is what the checking looks like 22(11) + 70= 17(11) - 95. I'm wondering did i make a mistake can you help?

hartnn (hartnn):

22x+70=17x-95 show your work, what did u do, we'll point out the error

OpenStudy (anonymous):

I think that's a better idea :)

OpenStudy (anonymous):

for checking i did 22(11) +70 = 17( 11)-95 v v 342 +70=187- 95 v v 312 =92 that's where the imbalance was and if you need the equation work tell me im unsure at this point

OpenStudy (anonymous):

?

hartnn (hartnn):

how did u get x=11, that was incorrect, we asked you to show steps on how you got 11

OpenStudy (anonymous):

x=-33, and this is correct i guess

OpenStudy (anonymous):

sorry i was confused. the equation work was 22x + 70 =17x -95 i subtracted 17x to both sides and gotten 15x and continued which was 15x+70=-95 i used my notes which said to now subtract 70 to -95 and it looked like this 15x +70= -95 -70=-70 and had gotten 165 and divided it by 15x and my answer was x=11

hartnn (hartnn):

22-17 = 5

hartnn (hartnn):

not 15

OpenStudy (anonymous):

You were on the right track just mixed up the numbers :)

OpenStudy (anonymous):

oh now i see yes i do that a lot i mix up my numbers all the time its a problem i have sorry thank you very much!!

hartnn (hartnn):

welcome ^_^ make sure you get x=-33

OpenStudy (anonymous):

i will i'm working out the problem right now as i type ^^ you all were a huge help thank you very much !

OpenStudy (anonymous):

Your welcome anytime! ;)